EQuestion Help A baskitball team sells tckets that cost $10, $20, o or VIP seats, $30 The teamhas sold 3102 tickots overal. It has solkd 182 more $20 tickets than $10 ickets The otal sales are $59 170 How many tckets of each kind have been sold? How many $10 tickets were sold?
Answer #1

Let no of $10 tickets be x
Then $20 tickets shall be x+182

$30 tickets =3102-(x+x+182)=2920-2x

Total price of $10=x*10=10x

$20=20(x+182)=20x+3640

$30= 30*(2920-2x)=87600-60x

Total revenue=10x+20x+3640+87600-60x=59170

x=1069

Hence there were 1069, $10 tickets that were sold